OpenStudy (anonymous):

Hello everyone. :) Can someone please help me solve this equation? Really confused & stuck! 2w/4 = 3+5w/6 The choices a,b,c, & d will be in the comments. THANKS!<3

OpenStudy (anonymous):

A. w = 9/11 B. w = 3/2 C. w = -3/2 D. w = -1/9 Please help me, ASAP. Thank you.